### v2

# How Focused Are LLMs? A Quantitative Study via Repetitive Deterministic Prediction Tasks

Abstract
We investigate the performance of large language models (LLMs) on repetitive deterministic prediction tasks and study how the sequence accuracy rate (SAR) scales with output length. Each such task involves the repetition of the same operation $n$ times. Examples of such tasks include letter replacement in letter strings following a given rule, integer addition, and multiplication of string operators in many-body quantum mechanics. If the LLM performs the task by a simple repetition algorithm, the success rate would follow an exponential decay with sequence length. In contrast, our experiments on leading LLMs reveal a sharp double-exponential drop beyond a characteristic length scale—an accuracy cliff marking a transition from reliable to unstable generation. This implies that LLMs fail to execute each operation independently. To explain this phenomenon, we propose a statistical-physics-inspired model capturing the competition between external conditioning from the prompt and internal interference among generated tokens. The model quantitatively reproduces the observed crossover and provides an interpretable link between attention-induced interference and sequence-level failure. Fitting the model to empirical results across multiple LLMs and tasks yields effective parameters that characterize the intrinsic error rate and error accumulation factor for each model-task pair, offering a principled framework for understanding the limits of deterministic accuracy in LLMs.
1 Introduction
Accuracy rate is among the most fundamental and widely adopted evaluation metrics in machine learning, especially for classification and other discriminative tasks. Its appeal lies in the clarity of its definition: each input has a well-defined ground-truth label, and the model’s prediction can be judged unambiguously as correct or incorrect. This binary criterion makes accuracy a natural measure of performance, enabling direct comparison across models and datasets.
However, this paradigm does not directly extend to large language models (LLMs). Many LLM applications, such as machine translation, summarization, or dialogue, are inherently sequence-to-sequence generation problems. In these settings, the model generates text probabilistically, token by token, sampling from a distribution of plausible continuations. Crucially, there is rarely a single “correct” output: multiple valid translations, summaries, or conversational responses may exist. Consequently, simple exact-match accuracy is insufficient for evaluation, and researchers rely instead on a suite of task-specific metrics (e.g., BLEU, ROUGE, perplexity, or human judgment) (papineni-etal-2002-bleu; lin-2004-rouge; liang2022holistic) that capture fluency, coherence, and relevance.
Yet this limitation is context-dependent. In scientific domains—including mathematics, physics, and computer science—the nature of the task often changes. Here, many problems admit a unique correct solution: a definite numerical value, a closed-form expression, or a precise code output. For instance, the question “what is 2342342+34423422342342+3442342” has only one acceptable answer, and the model’s response must exactly match the correct sequence of digits to be considered correct (cobbe2021training; hendrycks2021measuring; lewkowycz2022solving). In such applications, the probabilistic flexibility of LLMs does not diminish the importance of exact correctness. The evaluation once again reduces to a binary criterion: either the model outputs the correct sequence or it does not.
This observation motivates the introduction of a new evaluation concept tailored for LLMs in scientific applications: the Sequence Accuracy Rate (SAR). SAR measures the proportion of problem instances for which the LLM produces the exact correct sequence as output. By aligning the evaluation of LLMs with the well-established standard of accuracy in classification, SAR provides a rigorous and interpretable metric for assessing LLMs in domains where precision and correctness are paramount, as illustrated in Fig.˜1.
LLMs have achieved remarkable progress across a wide range of tasks, yet their reliability on structured, multi-step reasoning remains poorly understood (wei2022chain; havrilla2024glore). While in generative modeling the challenge has often been framed as a tradeoff between diversity and realism (astolfi2024consistency), multi-step reasoning tasks demand consistent correctness across sequential subtasks. Success in these settings is inherently factorizable: the probability of solving the overall problem is the product of per-step accuracies, making performance highly sensitive to individual step error rates. This exposes whether models can exploit compositional structure rather than treating complex problems monolithically (lake2018generalization).
To probe this challenge, we introduce a minimal and fully deterministic evaluation class where each problem instance admits a unique ground-truth solution and complexity can be tuned by a single parameter $N$. We present three complementary benchmarks: (i) cyclic letter replacement task, a purely symbolic problem over a finite alphabet that isolates stepwise error propagation without memory; (ii) integer addition, which tests the handling of local carry rules and long-range carry propagation across digits; and (iii) Pauli string multiplication, which requires tracking both local symbol rules and a global phase factor, thereby testing the model’s ability to maintain a lightweight memory across many steps All three benchmarks share the key property that overall complexity decays multiplicatively with sequence length $N$, providing a clean probe of cumulative error accumulation in deterministic reasoning (wang2022self; liu2023lost; turpin2023language; wan2025fano).
Figure 1: Illustration of the experimental setup and theoretical framework. The left panel shows examples of deterministic sequence prediction tasks (e.g., arithmetic problems) used to evaluate large language models (LLMs), where each token’s correctness is represented by a binary variable (Ising spin). Token errors are modeled as Ising spins interacting through all-to-all random couplings, capturing how correlations and noise propagate during sequence generation. The right panel displays the typical behavior of the Sequence Accuracy Rate (SAR) (i.e. the probability for all tokens to be correct in a output sequence) as a function of the sequence length $N$: SAR remains high for short sequences, then drops sharply beyond a characteristic crossover scale $N_{*}$, forming the accuracy cliff, which is nicely reproduced by a spin-glass statistical-mechanical model.
Beyond empirical evaluation, we develop a statistical physics interpretation of the observed behavior of SAR, drawing an analogy to the Ising model to assuming token correctness influence each other in random manner. Our theoretical model reproduces the accuracy-cliff cross-over behavior observed in LLMs, aligning closely with observed scaling behaviors, offering principled intuition for the limits of reliability as complexity grows. Together, our benchmarks and theoretical framework establish a controlled setting for studying how LLMs manage compositional reasoning, memory, and robustness in deterministic problem families.
2 Experiment
We prompted a set of large language models (LLMs)—gpt-5, gemini-2.5-pro, gemini-2.5-flash, grok-4, and claude-4-sonnet—to perform a series of deterministic sequence prediction tasks that each admit a unique correct solution. These tasks provide a well-defined setting for evaluating the sequence accuracy rate (SAR), enabling a quantitative analysis of how the accuracy of generated sequences scales with the problem size, measured by the output sequence length. During evaluation, all models were tested in a closed setting with external tools such as code execution and web access explicitly disabled, ensuring that results reflect the models’ internal reasoning capabilities only. While the quantitative results may vary slightly with prompt phrasing or instruction style, the qualitative behaviors reported here are robust across formulations. The prompt templates used for each task are available in the source code at GitHub.
2.1 Cyclic Letter Replacement
We start with a simple cyclic letter replacement task defined over an alphabet
$\mathcal{A}=\{\mathrm{A},\mathrm{B},\mathrm{C},\ldots\}.$ (1)
Each letter is mapped to an integer index $x\in\{0,1,\ldots,|\mathcal{A}|-1\}$. The transformation rule is a cyclic shift by one modulo the alphabet size:
$y=(x+1)\bmod|\mathcal{A}|.$ (2)
The output is then mapped back to a letter in $\mathcal{A}$. Each position in the input string is transformed independently, so for a string of length $N$ the task is simply to apply the shift rule to each site.
Formally, given an input string
$S=\big(x_{1},x_{2},\ldots,x_{N}\big),\quad x_{j}\in\{0,\ldots,|\mathcal{A}|-1\},$ (3)
the transformed output string is
$T=\big((x_{1}+1)\bmod|\mathcal{A}|,\ (x_{2}+1)\bmod|\mathcal{A}|,\ \ldots,\ (x_{N}+1)\bmod|\mathcal{A}|\big).$ (4)
Figure 2: Cyclic letter replacement benchmark on gemini-2.5-pro and gemini-2.5-flash with alphabet size $|\mathcal{A}|=4,9,13$ and $26$.
For example, for $\mathcal{A}={A,B,C,D}$:
Input: ADBAA $\Rightarrow$ Output: BACBB
Input: DABDC $\Rightarrow$ Output: ABCAD
For this task, the results are presented in Fig.˜2 and Fig.˜3. Fig.˜2 shows results from gemini-2.5-pro and gemini-2.5-flash with alphabet sizes $|\mathcal{A}|=4,9,13,26$. Fig.˜3 compares gpt-5, gemini-2.5-pro, gemini-2.5-flash, grok-4, and claude-4-sonnet at $|\mathcal{A}|=13$.
Figure 3: Cyclic letter replacement with alphabet size $|\mathcal{A}|=13$ across many different models.
2.2 Integer Addition
Next, we consider integer addition with a fixed number of digits. The task probes arithmetic consistency and multi-digit carry propagation—requiring sequential reasoning rather than purely single-site transformations.
Each query presents an addition problem of the form
$A+B,$ (5)
where $A,B\in\mathbb{N}$ are positive integers with the same digit lengths. The model is asked to compute the sum and return the result as a plain numeric string.
For example at $N=4$:
Input: $\texttt{1234}+\texttt{5678}$ $\Rightarrow$ Output: 6912
Input: $\texttt{9999}+\texttt{0001}$ $\Rightarrow$ Output: 10000
The corresponding results are shown in Fig.˜4, evaluating gemini-2.5-flash, gemini-2.5-pro, grok-4, and claude-4-sonnet.
Figure 4: Integer addition benchmark across different models.
2.3 Pauli Algebra
As a third deterministic benchmark, we consider Pauli string multiplication task with tunable string length. The Pauli operators $X$, $Y$, and $Z$ are $2\times 2$ matrices
$X=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right),~Y=\left(\begin{array}[]{cc}0&-i\\ i&0\end{array}\right),~Z=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right),$ (12)
They originate as generators of the Lie algebra $\mathfrak{su}(2)$, up to constants. For our purposes we treat them algebraically, using only their multiplication rules and phase factors. The identity operator is denoted $I$. The single-qubit multiplication rules are
$X\times X=Y\times Y=Z\times Z=I,$ (13)
$X\times Y=iZ,\quad Y\times X=-iZ,$ (14)
$Y\times Z=iX,\quad Z\times Y=-iX,$ (15)
$Z\times X=iY,\quad X\times Z=-iY,$ (16)
together with $I\times P=P$ for $P\in\{I,X,Y,Z\}$.
A Pauli string on $N$ sites is
$S=\alpha\bigotimes_{\ell=1}^{N}P_{\ell},$ (17)
where $\alpha\in\{\pm 1,\pm i\}$ is a global phase and $P_{\ell}\in\{I,X,Y,Z\}$. Because Pauli operators acting on different sites commute, the product of two strings reduces to independent local multiplications together with accumulation of the global phase:
$\left(\alpha_{1}\bigotimes_{\ell=1}^{N}P_{\ell}\right)\left(\alpha_{2}\bigotimes_{\ell=1}^{N}Q_{\ell}\right)=\Big(\alpha_{1}\alpha_{2}\prod_{\ell=1}^{N}\phi_{\ell}\Big)\bigotimes_{\ell=1}^{N}R_{\ell},$ (18)
where $P_{\ell}\times Q_{\ell}=\phi_{\ell}R_{\ell}$ with $\phi_{\ell}\in\{\pm 1,\pm i\}$ and $R_{\ell}\in\{I,X,Y,Z\}$.
This structure makes the task explicitly factorizable: each site requires a local Pauli multiplication and a phase update, while the overall result depends on consistent accumulation of all local outcomes. See App.˜A for a theoretical complexity analysis.
The need to consistently maintain and update the global phase highlights the role of memory handling in multi-step reasoning. Unlike purely sitewise problems, errors here can accumulate not only locally but also through the global phase register.
For this task, the results are presented in Fig.˜5, evaluating gemini-2.5-pro and gemini-2.5-flash under two evaluation criteria: one requiring both the operator sequence and accumulated phase to be correct, and a relaxed criterion considering only the operator sequence while ignoring phase errors.
Figure 5: Pauli string multiplication benchmark on gemini-2.5-pro and gemini-2.5-flash evaluated under strict (phase-matched) and relaxed (phase-ignored) criteria.
3 Analysis
3.1 Empirical Formula
In our experiments, we observed that the Sequence Accuracy Rate (SAR) depends strongly on the output sequence length $N$. Assuming independent per-token error events during generation, the baseline expectation is that the sequence accuracy rate decays exponentially with the sequence length $N$:
$\text{SAR}(N)=\mathrm{e}^{-\beta N},$ (19)
where $\beta>0$ denotes a constant per-token error rate. However, as we have seen in the above experiments, the empirical results deviate significantly from this simple picture. Instead of a smooth exponential decay, SAR exhibits a distinct crossover behavior: it remains nearly perfect for short sequences and then undergoes a rapid drop to zero around a characteristic length scale $N_{*}$.
This observation indicates that token errors are not independent: earlier tokens affect the error rate of subsequent tokens. To capture this history dependence, we treat the per-token error rate as a function of the sequence length $N$ rather than a constant and adopt the following empirical fit:
$\text{SAR}(N)=\mathrm{e}^{-\beta(N)N},\quad\beta(N)=\beta_{0}\,\alpha^{N-1},$ (20)
where $\beta_{0}>0$ quantifies the intrinsic error rate, and $\alpha$ represents the error accumulation factor, which characterizes how local errors compound as the sequence length increases. When $\alpha=1$, the model reduces to the standard exponential decay expected from independent token errors. For $\alpha>1$, however, the effective error rate $\beta(N)$ grows exponentially with $N$ even when $\beta_{0}$ is small, leading to a rapid, catastrophic drop in accuracy once $\beta(N)$ becomes of order one. The onset of this accuracy cliff occurs at a characteristic sequence length
$N_{*}\simeq 1+\frac{\log(1/\beta_{0})}{\log\alpha},$ (21)
beyond which scale the model’s reliability collapses.
3.2 Hypothesis and Statistical Model
Empirically, the observed crossover length $N_{*}$ is typically on the order of $10\sim 100$—well within the attention window ($\sim 10^{5}$ or larger) of modern LLMs. This rules out limited context length as the primary cause of the accuracy cliff. We therefore hypothesize that the rapid accuracy degradation arises intrinsically from the all-to-all coupling structure of the self-attention mechanism. While the attention weights in next-token prediction are structured and typically sparse rather than uniformly distributed over all past tokens, the architecture nonetheless enables fluctuations and errors to spread through the network in an effectively all-to-all fashion, propagating globally across the entire generated sequence. As a result, small local errors can be amplified and spread through these dense interactions, effectively inducing collective fluctuations among token-level accuracies.
To analyze this mechanism, we abstract away token content and focus solely on whether each token is generated correctly. We introduce binary variables $s_{i}=\pm 1$ to represent the correctness of the $i$-th token, with $s_{i}=+1$ for correct and $s_{i}=-1$ for incorrect generation. We model the joint distribution of token errors $s=\{s_{i}\}$ using an effective Ising-like energy function,
$E_{J}[s]=-\sum_{i<j}J_{ij}\,s_{i}s_{j}-h\sum_{i}s_{i},$ (22)
where $h$ denotes an external field reflecting the baseline bias from the prompt toward correct generation, and $J_{ij}$ represent effective couplings between token correctness variables. In the independent-token limit, $J_{ij}=0$ and the model will predict the exponential decay SAR following Eq. (19). However, experimental results described by Eq. (20) clearly indicate deviations from this independent picture, implying the presence of correlations among token outcomes.
We attribute these correlations to noisy, attention-induced interactions that can be modeled as random couplings drawn independently from a Gaussian ensemble,
$\mathbb{E}_{J}[J_{ij}]=0,\qquad\mathbb{E}_{J}[J_{ij}^{2}]=J_{0}^{2},$ (23)
where $J_{0}$ characterizes the strength of internal interference or “noise coupling.” The random couplings capture how attention-mediated dependencies introduce both constructive and destructive interference among tokens. Together with the external field h, which sets the intrinsic per-token accuracy, this formulation defines a statistical-physics model for the collective behavior of token errors. The model is mathematically equivalent to the Sherrington–Kirkpatrick (SK) spin-glass model, widely studied in statistical physics.
For a given realization of the couplings $J_{ij}$, the probability of observing a sequence of token correctness variables $s$ is
$p_{J}[s]\;=\;\frac{1}{Z_{J}}\,e^{-E_{J}[s]},$ (24)
with the partition function
$Z_{J}=\sum_{s}e^{-E_{J}[s]}.$ (25)
In particular, the SAR is associated with the probability of generating the fully correct sequence $s=(1,1,\ldots,1)$, averaged over disorder realizations of the couplings. Using the geometric mean to account for multiplicative scaling, we define
$\mathrm{SAR}\;:=\;\exp\!\Big(\mathbb{E}_{J}\big[\log p_{J}[s=1]\big]\Big).$ (26)
Under a perturbative approximation App.˜B, the solution of Eq.˜26 takes the same form as the empirical law in Eq.˜20, with empirical parameters $\alpha,\beta_{0}$ related to the energy model parameters $J_{0},h$ as follows:
$\alpha\;\simeq\;\mathrm{e}^{\frac{2J_{0}^{2}}{1+2J_{0}}},\qquad\beta_{0}\;=\;\mathrm{e}^{-2h}.$ (27)
From the statistical-physics perspective, the accuracy cliff corresponds to a finite-size crossover in the underlying spin system. For small system size $N$, the external field $h$ dominates, and the system behaves as a paramagnet, aligning most spins in the correct direction ($s_{i}=+1$). As $N$ increases, however, the collective spin correlation energy from the random couplings $J_{ij}$, which scales as $N^{2}$, eventually exceeds the entropic contribution that grows only linearly with $N$. Beyond a characteristic size $N_{*}$, the interaction energy dominates, and the system crosses over into a spin-glass-like regime, where the fully aligned configuration is no longer favored, leading to the rapid accuracy degradation observed as the accuracy cliff.
Using the fitted results of the numerical data based on the form in Eq.˜20, we pin each model onto the correlation–error map. This mapping translates observed SAR decay profiles into quantitative measures of intrinsic error rate and correlation strength, enabling direct comparison of how different LLMs accumulate and propagate errors across tasks. Each point on the map represents a model’s performance under a specific task, corresponding to a pair $(\log\alpha,\log\beta_{0})$ with color indicating $\log N_{*}$, where $N_{*}$ is defined as the solution of $\mathrm{SAR}(N_{*})=0.5$ under the given parameters. A smaller $\log\alpha$ means the model has a sharper focus of attention, and a smaller $\log\beta_{0}$ means the model has a lower error rate for executing a single unit task. The mapping results grouped by model are shown in Fig.˜6, and those grouped by task are shown in Fig.˜7.
Figure 6: Correlation–error map grouped by model, with color indicating the value of $\log N_{*}$. The horizontal and vertical axes, $(\log\alpha,\log\beta_{0})$, represent the correlation level and error level, respectively. The gray contour line and the numbers along it denote the sequence length $N$.
Figure 7: Correlation–error map grouped by task, with color indicating the value of $\log N_{*}$. The horizontal and vertical axes, $(\log\alpha,\log\beta_{0})$, represent the correlation level and error level, respectively. The gray contour line and the numbers along it denote the sequence length $N$.
4 Implication: Divide-and-Conquer Strategy
The empirical scaling law of Eq. (20) not only provides a qualitative description of how the accuracy inevitably degrades with the output sequence length $N$ for LLMs, but also offers a constructive implication: when $\alpha>1$, the reliability of long-sequence generation can be improved by adopting a divide-and-conquer strategy.
In this context, divide-and-conquer refers to partitioning a sequence-generation task of total length $N$ into $k$ sub-tasks of length $N/k$, prompting the model separately for each sub-task, and then merging the partial outputs to form the complete sequence. The intuition is that the observed accuracy cliff originates from correlated error propagation among tokens: when $\alpha>1$, the effective per-token error rate grows exponentially with $N$, driven by all-to-all dependencies in the attention mechanism. Dividing the sequence into shorter fragments effectively cuts the correlation loops, preventing the catastrophic error accumulation.
We introduce a multiplicative overhead factor $\theta_{N,k}\in(0,1]$ to account for the reduction of SAR due to the additional operations involved in applying the divide-and-conquer strategy, such as segmentation, formatting, boundary handling, and content copying. $\theta_{N,k}$ is expected to only depend on $N$ and $k$ weakly. The overall success probability under divide-and-conquer is then
$\mathrm{SAR}_{\mathrm{DC}}(N,k)=\theta_{N,k}\,\big[\mathrm{SAR}(N/k)\big]^{k}=\theta_{N,k}\,\exp\!\Big[-\beta_{0}\,N\,\alpha^{\,\frac{N}{k}-1}\Big].$ (28)
Define the logarithmic gain of divide-and-conquer relative to the single-shot strategy as
$\Delta(N,k):=\log\frac{\mathrm{SAR}_{\mathrm{DC}}(N,k)}{\mathrm{SAR}(N)}=\log\theta_{N,k}+\beta_{0}N\Big(\alpha^{\,N-1}-\alpha^{\,\frac{N}{k}-1}\Big).$ (29)
This formulation shows that for $\alpha>1$, the gain term increases rapidly with $N$, indicating that the larger $\alpha$ is, the more advantageous segmentation becomes.
Theorem 1 (Advantage of Divide-and-Conquer).
Let the intrinsic error rate $\beta_{0}>0$ and the correlation amplification factor $\alpha>1$, and fix the integer $k\geq 2$. Under the empirical scaling law $\mathrm{SAR}(N)=\exp[-\beta_{0}N\alpha^{\,N-1}]$, for the segmentation into $k$ equal sub-tasks to yield a positive gain $\Delta(N,k)>0$, it would be sufficient if the sequence length exceeds the following bound
$N\geq N_{\mathrm{DC}}\;=\;1+\frac{1}{\log\alpha}\left(\log\left(1-\frac{2\log\theta_{N,k}}{\beta_{0}}\right)+\frac{\log 2}{1-1/k}\right),$ (30)
meaning that for all $N\geq N_{\mathrm{DC}}$, one has $\mathrm{SAR}_{\mathrm{DC}}(N,k)>\mathrm{SAR}(N)$.
(See App. C for a proof.)
This result provides an explicit criterion for when the divide-and-conquer approach is beneficial. Under this strategy, as evident from Eq. (28), the intrinsic error rate $\beta_{0}$ remains unchanged, while the correlation amplification factor $\alpha$ is effectively rescaled to $\alpha_{\text{eff}}\simeq\alpha^{1/k}$ in the large-$N$ regime. As a consequence, the characteristic accuracy-cliff scale is extended to
$N_{*}\simeq 1+\frac{\log(1/\beta_{0})}{\log\alpha_{\text{eff}}}=1+k\,\frac{\log(1/\beta_{0})}{\log\alpha},$ (31)
growing approximately linearly with $k$, according to Eq. (21). The empirical effectiveness of the divide-and-conquer strategy is demonstrated in Fig.˜8, where the linear-$k$ scaling of accuracy cliff is indeed observed.
In summary, the empirical scaling law does more than describe the accuracy cliff—it also predicts how to mitigate it: segmentation effectively reduces the compounding of correlated errors and extends the reliable length scale of LLMs by roughly a factor of $k$, transforming an observed limitation into a quantitative improvement strategy.
Figure 8: Divide-and-conquer evaluation on the Pauli string multiplication task under both phase-restricted and phase-relaxed criteria, using gemini-2.5-flash and gemini-2.5-pro. The gray k=1k{=}1 curves show the fitted $\mathrm{SAR}(N)$ from Eq.˜20; the extracted $(\alpha,\beta_{0})$ values are then used to plot the theoretical $\mathrm{SAR}(N/k)^{k}$ curves for k=2k{=}2 and k=3k{=}3 (red and blue), assuming no overhead ($\theta_{N,k}=1$). In the phase-restricted case, the pro model closely follows the predicted improvement, while the flash model performs below the theoretical curves, due to additional overhead in divide and combine. In the phase-relaxed case, both models closely follow the predicted improvement, and additionally expand the perfect-accuracy plateau by approximately a factor of $k$ (indicated by the colored dashed lines).
5 Related Works
Single-pass reasoning and capacity limits. Recent research has shown that even when models can locally apply correct reasoning steps, global accuracy collapses as reasoning chains lengthen. This phenomenon parallels information-theoretic capacity limits of single forward passes. The Fano-style upper bound framework  (wan2025fano) formalizes this collapse for multi-hop question answering (MHQA), proving that once task complexity exceeds the model’s per-pass information capacity, accuracy must drop sharply—mirroring our empirical SAR crossover. Their work interprets the reasoning chain as a noisy communication channel, bounded by mutual information, while our SAR analysis arrives at the same breakdown curve from a statistical-physics viewpoint, emphasizing the all-to-all error propagation in the attention architecture as the key underlying mechanism. Together they highlight the inevitability of reliability loss in single-pass reasoning under finite model capacity.
Prompting paradigms and decomposition. A wide range of prompting methods—Direct prompting, Chain-of-Thought (CoT) (wei2022chain; kojima2022large), ReAct (yao2023react), Plan-and-Solve (wang2023plan), and Self-Ask (press2023measuring)—operate in the single-pass regime, assuming the full reasoning chain fits within one generation. By contrast, multi-call frameworks such as Self-Refine (madaan2023self), InfoQA (wan2025fano), and agentic systems for programming or reasoning (li2024survey; qian2024chatdev; kim2024llm) mitigate this bottleneck by decomposing reasoning into smaller, verifiable subcalls. Our SAR-based analysis provides a quantitative explanation of why such multi-call methods succeed: by resetting the error accumulation process, they effectively bound per-step complexity below the model’s capacity threshold, preventing the exponential or super-exponential degradation seen in single-pass reasoning.
Self-checking, verification, and robust reasoning. Other orthogonal strategies—such as self-consistency sampling, debate-style verification, and tool-augmented reasoning—attempt to mitigate compounding errors by introducing redundancy or external validation (yao2023react; xie2025fire; wan2025cognitive; mccoy2023embers). SAR offers a unified diagnostic lens for such methods: improved reliability corresponds to reduced intrinsic error rate $\beta_{0}$, while better verification or decomposition corresponds to smaller accumulation factor $\alpha$. Thus, SAR quantitatively links algorithmic interventions to empirical stability regimes.
6 Conclusion and Future Work
We introduced the Sequence Accuracy Rate (SAR) as a principled metric for quantifying reliability in deterministic reasoning tasks performed by large language models. Through three controlled benchmarks—cyclic letter replacement, integer addition, and Pauli string multiplication—we demonstrated that SAR provides a quantitative view of how local inaccuracies compound with sequence length. Empirical results reveal a sharp reliability crossover rather than a simple exponential decay, indicating correlated error accumulation rather than independent token noise.
To explain this behavior, we developed a statistical-physics model inspired by the Sherrington–Kirkpatrick spin-glass system, linking empirical scaling parameters to interpretable quantities: intrinsic error rate $\beta_{0}$ and error accumulation strength $\alpha$. This hypothetical model connects observed accuracy cliffs in deterministic reasoning to correlated interactions within LLMs.
Future directions include extending SAR to stochastic reasoning pipelines, integrating it with self-verification or multi-call decomposition schemes, and applying the statistical-physics approach to study model scaling laws and architectural modifications. More broadly, SAR provides a foundation for treating reasoning reliability as a measurable and scalable property, enabling systematic analysis and engineering of models that sustain compositional accuracy over increasing complexity.
7 Acknowledgment
We would like to thank Chen Nie for the engineering support. We would like to acknowledge helpful discussion with Victor Albert and Hsin-Yuan Huang on bench.science. W.H. additionally thanks Tianxiao Hu for helpful discussions.
Appendix A Success Rate Analysis of Pauli String test
Suppose an intellectual agent attempts to compute the product of two Pauli strings of length $N$. At each site, two types of operations are required:
  1. 1.
    A Pauli multiplication step, where $P_{i}\times Q_{i}$ is reduced to a Pauli operator $R_{i}$ and a local phase factor $\phi_{i}$. Let the probability of correctly performing this step be $p_{\sigma}$.
  2. 2.
    A phase multiplication step, where the accumulated phase is updated as $\Phi\mapsto\Phi\times\phi_{i}$. Let the probability of correctly performing a single phase update be $p_{\phi}$.
We first model the accumulated phase as a Markov chain on the cyclic group $G=\{1,i,-1,-i\}$ with the state order (1,i,1,i)(1,i,-1,-i). At each step, the algorithm multiplies the current phase by one of the group elements, but due to noise it returns an incorrect element with probability $p_{\phi}$, and the correct (true) element with probability $1-p_{\phi}$. Hence, each of the three incorrect elements occurs with probability $\tfrac{p_{\phi}}{3}$. Let $\mathds{1}$ denote the $4\times 4$ identity matrix and $J$ the all-ones matrix. When the intended multiplier is $1$, the one-step transition kernel is the circulant matrix
$K=(1-p_{\phi})\,\mathds{1}+\frac{p_{\phi}}{3}\,(J-\mathds{1}).$ (32)
Define the four permutation (left-shift) matrices $\{S_{\phi}:\phi\in G\}$ that implement left multiplication by $\phi$ on the state order; explicitly,
$S_{1}=\mathds{1},\qquad S_{i}=\begin{bmatrix}0&0&0&1\\ 1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix},\qquad S_{-1}=\begin{bmatrix}0&0&1&0\\ 0&0&0&1\\ 1&0&0&0\\ 0&1&0&0\end{bmatrix},\qquad S_{-i}=\begin{bmatrix}0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ 1&0&0&0\end{bmatrix}$.
Thus, if the multiplier at a step is $\phi\in G$, the corresponding one-step transition matrix is
$T(\phi)\;=\;S_{\phi}\,K.$ (33)
Because every permutation matrix commutes with $K$, i.e. $S_{\phi}K=KS_{\phi},\forall\phi\in G$, the product of step-wise transitions reduces to a similarity transform. For any sequence of true multipliers $\phi_{1},\ldots,\phi_{N}$, we have
$T(\phi_{N})\cdots T(\phi_{1})=\big(S_{\phi_{N}}\cdots S_{\phi_{1}}\big)\,K^{N}=S_{\Phi}\,K^{N},\quad\Phi:=\phi_{N}\cdots\phi_{1}.$ (34)
Therefore, the success rate of phase computing task is given by:
$P_{\phi}(N)=e_{\Phi}^{\top}S_{\Phi}\,K^{N}e_{1},$ (35)
where $e_{\Phi}$ is the basis vector of $\Phi$ in the state order (1,i,1,i)(1,i,-1,-i) and $K^{N}=\tfrac{1}{4}J+\Big(\tfrac{3-4p_{\phi}}{3}\Big)^{N}(\mathds{1}-\tfrac{1}{4}J)$. Explicitly, this gives:
$P_{\phi}(N)=\tfrac{1}{4}+\tfrac{3}{4}\Big(\tfrac{3-4p_{\phi}}{3}\Big)^{N}.$ (36)
Combining this with the individual Pauli multiplication success rate, the overall sequence accuracy rate is
$\text{SAR}(N)=(1-p_{\sigma})^{N}\left[\tfrac{1}{4}+\tfrac{3}{4}\left(\tfrac{3-4p_{\phi}}{3}\right)^{N}\right].$ (37)
The requirement to accumulate local phase factors across all sites is especially instructive, as it directly reflects the need for memory handling in agent design: just as an LLM must correctly maintain and update a state variable across multiple reasoning steps, the phase register here must be updated consistently at every site.
Appendix B Modeling the LLM SAR with SK-Model
In this appendix, we propose and solve a minimal energy-based toy model that captures the empirical scaling behavior in Eq.˜20 and offers a possible physical interpretation of its underlying mechanism.
Let $\mathcal{D}=\{(x_{(k)},y_{(k)})\}_{k=1}^{K}$ be a dataset consisting of input–output sequence pairs, where each input sequence $x_{(k)}$ is deterministically mapped to a unique correct output sequence $y_{(k)}$. An LLM trained on this dataset defines a conditional probability distribution $p_{\theta}(y\mid x)$ over output sequences.
B.1 Definitions and independent-token baseline
The Sequence Accuracy Rate (SAR) is defined as the geometric mean of the model’s probability on correct sequences:
$\mathrm{SAR}(p_{\theta};\mathcal{D})=\exp\!\Big(\frac{1}{K}\sum_{k=1}^{K}\log p_{\theta}(y_{(k)}\mid x_{(k)})\Big)\in(0,1].$ (38)
Introduce Ising variables for correctness:
$s_{i}=\begin{cases}+1,&y_{i}\ \text{is correct},\\ -1,&\text{otherwise},\end{cases}\qquad s=(s_{1},\dots,s_{N})\in\{\pm 1\}^{N},$
so all-correct output sequence corresponds to $s_{i}=+1$ for all $i=1,\ldots,N$, which we denote jointly as $s=\mathbf{1}$. If each token were generated independently with a token-level accuracy rate $p_{\mathrm{tok}}$, the expected sequence accuracy would scale as
$\mathrm{SAR}=p_{\mathrm{tok}}^{N},$ (39)
which decays exponentially with sequence length $N$. The corresponding energy model for the Ising variables is simply
$E[s]=-h\sum_{i=1}^{N}s_{i},$ (40)
such that
$\mathrm{SAR}=\frac{e^{-E[s=\mathbf{1}]}}{\sum_{s}e^{-E[s]}}=\left(\frac{1}{1+e^{-2h}}\right)^{N},$ (41)
implying that the external field $h$ parameterizes the token-level accuracy as
$p_{\mathrm{tok}}=\frac{1}{1+e^{-2h}}.$ (42)
However, this independent-token prediction fails to capture the crossover behavior observed in experiments.
B.2 Energy model (Sherrington–Kirkpatrick (SK) Hamiltonian)
In our experiments, the Sequence Accuracy Rate (SAR) shows a pronounced dependence on output length $N$. While one might expect a simple exponential decay, $\mathrm{SAR}\sim e^{-\alpha N}$, our results reveal a clear accuracy cliff: SAR remains high for short sequences, then sharply declines around a characteristic length. This behavior indicates that sequence-level reliability in LLMs arises not from independent token errors but from correlated or scale-dependent effects, such as contextual coupling and calibration across sequence lengths, governing the shift from reliable to unreliable generation in deterministic tasks. We hypothesize that the deviation arises from all-to-all correlations among tokens during sequence generation. The self-attention mechanism in LLMs inherently couples every token to every other token, introducing dependencies beyond the independent-token picture. These correlations are not purely constructive: noise, interference, or internal competition within the network can generate effective random coupling strengths $J_{ij}$ between the Ising variables $s_{i}$ (the token correctness labels). Such couplings modify the collective statistics of sequence correctness, producing a non-exponential decay and the observed crossover in SAR.
We therefore propose the following energy model:
$E_{J}[s]=-\sum_{i<j}J_{ij}\,s_{i}s_{j}-h\sum_{i=1}^{N}s_{i},$ (43)
where $s_{i}=\pm 1$ for $i=1,2,\ldots,N$ are Ising variables representing the correctness of the $i$-th token $y_{i}$ in the output sequence generated by the LLM.
The coupling strengths $J_{ij}$ are drawn independently from a Gaussian distribution with zero mean and finite variance:
$\mathbb{E}_{J}[J_{ij}]=0,\qquad\mathbb{E}_{J}[J_{ij}^{2}]=J_{0}^{2}.$ (44)
They model the noisy all-to-all correlations in token generation, and $J_{0}$ characterizes the noise strength, which is assumed to be weak. The external field $h$ parameterizes the token-level accuracy rate via Eq. (0): a strong field $h$ corresponds to a high token-level accuracy, which is typically assumed. This model is equivalent to the well-known Sherrington–Kirkpatrick (SK) model in spin-glass physics.
For a particular random realization of $J_{ij}$, the probability of observing the Ising sequence $s$ is
$p_{J}[s]=\frac{1}{Z_{J}}\,e^{-E_{J}[s]},$ (45)
where $Z_{J}$ is the partition function,
$Z_{J}=\sum_{s}e^{-E_{J}[s]}.$ (46)
Specifically, the Sequence Accuracy Rate (SAR) corresponds to the geometric mean probability of the all-correct configuration under random couplings:
$\mathrm{SAR}:=\exp\!\big(\mathbb{E}_{J}\log p_{J}[s=\mathbf{1}]\big).$ (47)
B.3 Replica trick and disorder average
From Eq.˜47, it is straightforward to show that
$\log\mathrm{SAR}$ $=\mathbb{E}_{J}\log p_{J}[s=\mathbf{1}]=\mathbb{E}_{J}\log\!\left(\frac{e^{-E_{J}[s=\mathbf{1}]}}{Z_{J}}\right)$
$=\mathbb{E}_{J}\big(-E_{J}[s=\mathbf{1}]-\log Z_{J}\big)=\mathbb{E}_{J}\!\Big(\sum_{i<j}J_{ij}+hN-\log Z_{J}\Big).$ (48)
Using the disorder mean $\mathbb{E}_{J}[J_{ij}]=0$, we have
$\log\mathrm{SAR}=hN-\mathbb{E}_{J}\log Z_{J}.$ (49)
Replica trick.
The expectation $\mathbb{E}_{J}\log Z_{J}$ in (49) is intractable directly. Physicists therefore use the replica trick:
$\mathbb{E}_{J}\log Z_{J}=\lim_{n\to 0}\partial_{n}\,\mathbb{E}_{J}\big[Z_{J}^{\,n}\big].$ (50)
By definition,
$\mathbb{E}_{J}\big[Z_{J}^{\,n}\big]$ $=\mathbb{E}_{J}\left[\Big(\sum_{[s]}\exp\big(\sum_{i<j}J_{ij}s_{i}s_{j}+h\sum_{i}s_{i}\big)\Big)^{\!n}\right]$
$=\sum_{\{s^{a}\}}\exp\!\Bigg(\sum_{a=1}^{n}\sum_{i<j}J_{ij}\,s_{i}^{a}s_{j}^{a}+h\sum_{a=1}^{n}\sum_{i=1}^{N}s_{i}^{a}\Bigg),$ (51)
where the spin variables have been replicated to $s_{i}^{a}$ with $i=1,\ldots,N$ and $a=1,\ldots,n$. Here $N$ is the number of sites (sequence length) and $n$ the number of replicas.
Disorder average.
Averaging $Z_{J}^{\,n}$ over i.i.d. Gaussian couplings $J_{ij}$ with $\mathbb{E}_{J}[J_{ij}]=0$, $\mathbb{E}_{J}[J_{ij}^{2}]=J_{0}^{2}$, we obtain
$\mathbb{E}_{J}\big[Z_{J}^{\,n}\big]$ $=\sum_{\{s^{a}\}}\prod_{i<j}\mathbb{E}_{J_{ij}}\exp\!\Big(\sum_{a=1}^{n}J_{ij}\,s_{i}^{a}s_{j}^{a}\Big)\,\exp\!\Big(h\sum_{a=1}^{n}\sum_{i=1}^{N}s_{i}^{a}\Big)$
$=\sum_{\{s^{a}\}}\prod_{i<j}\exp\!\Big(\frac{J_{0}^{2}}{2}\,C_{ij}^{2}\Big)\,\exp\!\Big(h\sum_{a=1}^{n}\sum_{i=1}^{N}s_{i}^{a}\Big),$ (52)
where $C_{ij}\equiv\sum_{a=1}^{n}s_{i}^{a}s_{j}^{a}$ and we used the Gaussian moment-generating function $\mathbb{E}_{J_{ij}}e^{J_{ij}C_{ij}}=\exp\!\big(\tfrac{J_{0}^{2}}{2}C_{ij}^{2}\big).$ Thus,
$\mathbb{E}_{J}\big[Z_{J}^{\,n}\big]$ $=\sum_{\{s^{a}\}}\exp\!\Big(\frac{J_{0}^{2}}{2}\sum_{i<j}C_{ij}^{2}+h\sum_{a=1}^{n}\sum_{i=1}^{N}s_{i}^{a}\Big)$
$=\sum_{\{s^{a}\}}\exp\!\Big(\frac{J_{0}^{2}}{2}\sum_{i<j}\big(\sum_{a=1}^{n}s_{i}^{a}s_{j}^{a}\big)^{2}+h\sum_{a=1}^{n}\sum_{i=1}^{N}s_{i}^{a}\Big).$ (53)
Expanding the quadratic replica term gives
$\sum_{i<j}\Big(\sum_{a=1}^{n}s_{i}^{a}s_{j}^{a}\Big)^{2}$ $=\sum_{i<j}\left(\sum_{a=1}^{n}(s_{i}^{a}s_{j}^{a})^{2}+2\sum_{a<b}s_{i}^{a}s_{j}^{a}s_{i}^{b}s_{j}^{b}\right)$
$=\sum_{i<j}\left(n+2\sum_{a<b}s_{i}^{a}s_{j}^{a}s_{i}^{b}s_{j}^{b}\right)$
$=\frac{nN(N-1)}{2}+2\sum_{i<j}\sum_{a<b}s_{i}^{a}s_{j}^{a}s_{i}^{b}s_{j}^{b}.$ (54)
Introducing the constant $\Omega\equiv\exp\!\big(J_{0}^{2}N(N-1)/4\big)$ to absorb the $\mathcal{O}(n)$ term, we obtain the compact form
$\mathbb{E}_{J}\big[Z_{J}^{\,n}\big]=\Omega^{\,n}\sum_{\{s^{a}\}}\exp\!\Big(J_{0}^{2}\sum_{i<j}\sum_{a<b}s_{i}^{a}s_{j}^{a}s_{i}^{b}s_{j}^{b}+h\sum_{a=1}^{n}\sum_{i=1}^{N}s_{i}^{a}\Big).$ (55)
Geometric interpretation.
Equation (55) defines a plaquette four-spin interaction on the graph product $K_{N}\times K_{n}$ (complete graphs on $N$ and $n$ vertices, respectively). The construction is self-dual under the exchange $N\leftrightarrow n$.
Figure 9: Example: $K_{4}\times K_{3}$ graph.
B.3.1 Small-$J_{0}$ Expansion
Assuming that the noise strength $J_{0}$ is weak, we perform a perturbative expansion of the replicated partition function in powers of $J_{0}$:
$\mathbb{E}_{J}Z_{J}^{\,n}$ $=\Omega^{n}\sum_{\{s\}}e^{\,h\sum_{i,a}s_{i}^{a}}\Bigg[1+J_{0}^{2}\sum_{i<j,a<b}s_{i}^{a}s_{j}^{a}s_{i}^{b}s_{j}^{b}+\frac{J_{0}^{4}}{2}\Bigg(\sum_{i<j,a<b}s_{i}^{a}s_{j}^{a}s_{i}^{b}s_{j}^{b}\Bigg)^{\!2}+\cdots\Bigg].$ (56)
Because the Boltzmann factor factorizes over spins,
$e^{h\sum_{i,a}s_{i}^{a}}=\prod_{i,a}e^{hs_{i}^{a}},$
the spin summations can be carried out independently. For a single spin, we define
$z:=\sum_{s=\pm 1}e^{hs}=2\cosh h,\qquad m:=\langle s\rangle=\frac{1}{z}\sum_{s=\pm 1}s\,e^{hs}=\tanh h.$ (57)
Hence, after performing the average over independent spins,
$\mathbb{E}_{J}Z_{J}^{\,n}$ $=\Omega^{n}z^{nN}\Bigg[1+\frac{J_{0}^{2}}{4}N(N-1)n(n-1)m^{4}$
$\quad+\frac{J_{0}^{4}}{2}N(N-1)n(n-1)\Big(\tfrac{1}{4}+\tfrac{1}{2}(N+n-4)m^{4}+(N-2)(n-2)m^{6}+$ (58)
$\tfrac{1}{4}\big(\tfrac{1}{4}N(N-1)n(n-1)-(2N-3)(2n-3)\big)m^{8}\Big)+\cdots\Bigg].$ (59)
Plaquette pairings.
The term
$\big\langle\sum_{i<j,a<b}s_{i}^{a}s_{j}^{a}s_{i}^{b}s_{j}^{b}\big\rangle=\tfrac{1}{4}N(N-1)n(n-1)m^{4}$
counts the number of plaquettes on the product graph $K_{N}\times K_{n}$, corresponding to choosing one edge from $K_{N}$ and one from $K_{n}$ independently. Each summand represents a single four-spin plaquette interaction.
At order $J_{0}^{4}$, we must evaluate $\big\langle\big(\sum_{i<j,a<b}s_{i}^{a}s_{j}^{a}s_{i}^{b}s_{j}^{b}\big)^{2}\big\rangle$, which decomposes into five distinct topological classes of plaquette pairings:
  • (A)
    Double-covered plaquettes: $\tfrac{1}{4}N(N-1)n(n-1)$.
  • (B)
    Edge-sharing plaquette pairs (sharing a $K_{n}$ edge): $\tfrac{1}{2}N(N-1)(N-2)n(n-1)m^{4}$.
  • (B)
    Edge-sharing plaquette pairs (sharing a $K_{N}$ edge): $\tfrac{1}{2}N(N-1)n(n-1)(n-2)m^{4}$.
  • (C)
    Corner-sharing plaquette pairs: $N(N-1)(N-2)n(n-1)(n-2)m^{6}$.
  • (D)
    Non-intersecting plaquette pairs: $\tfrac{1}{4}N(N-1)n(n-1)\big[\tfrac{1}{4}N(N-1)n(n-1)-(2N-3)(2n-3)\big]m^{8}$.
The sum of all classes yields
$\Big\langle\Big(\sum_{i<j,a<b}s_{i}^{a}s_{j}^{a}s_{i}^{b}s_{j}^{b}\Big)^{2}\Big\rangle$ $=N(N-1)n(n-1)\Big[\tfrac{1}{4}+\tfrac{1}{2}(N+n-4)m^{4}+(N-2)(n-2)m^{6}$
$\qquad\qquad+\tfrac{1}{4}\big(\tfrac{1}{4}N(N-1)n(n-1)-(2N-3)(2n-3)\big)m^{8}\Big].$ (60)
Substituting (60) into (59) gives the final expansion
$\mathbb{E}_{J}Z_{J}^{\,n}$ $=\Omega^{n}(2\cosh h)^{nN}\Big[1+\tfrac{J_{0}^{2}}{4}N(N-1)n(n-1)m^{4}$
$\quad+\tfrac{J_{0}^{4}}{2}N(N-1)n(n-1)\Big(\tfrac{1}{4}+\tfrac{1}{2}(N+n-4)m^{4}+(N-2)(n-2)m^{6}+$ (61)
$\tfrac{1}{4}\big(\tfrac{1}{4}N(N-1)n(n-1)-(2N-3)(2n-3)\big)m^{8}\Big)+\cdots\Big].$ (62)
Replica limit.
Taking the replica limit as in Eq. (50),
$\mathbb{E}_{J}\log Z_{J}$ $=\lim_{n\to 0}\partial_{n}\,\mathbb{E}_{J}Z_{J}^{\,n}$
$=\log\Omega+N\log(2\cosh h)-\frac{J_{0}^{2}}{4}N(N-1)m^{4}$
$\quad-\frac{J_{0}^{4}}{2}N(N-1)\Big[\tfrac{1}{4}+\tfrac{1}{2}(N-4)m^{4}-2(N-2)m^{6}+\tfrac{3}{4}(2N-3)m^{8}\Big]+\mathcal{O}(J_{0}^{6}).$ (63)
Result for SAR.
Substituting (63) into Eq. (49) yields
$\log\mathrm{SAR}=N\big(h-\log(2\cosh h)\big)+\frac{J_{0}^{2}}{4}N(N-1)\big(-1+\tanh^{4}h\big)$
$\quad+\frac{J_{0}^{4}}{2}N(N-1)\Big[-\tfrac{1}{2}-(N-4)\tanh^{4}h+4(N-2)\tanh^{6}h-\tfrac{3}{2}(2N-3)\tanh^{8}h\Big]+\mathcal{O}(J_{0}^{6}).$ (64)
For $h\gg 1$, using $\tanh h\approx 1-2e^{-2h}$,
$\log\mathrm{SAR}\simeq-e^{-2h}N\Big[\,1+2J_{0}^{2}(N-1)+2J_{0}^{4}(N-1)^{2}+\mathcal{O}(J_{0}^{6})\,\Big].$ (65)
Motivated by the expansion above, an empirical expression that reproduces the observed crossover behavior is
$\mathrm{SAR}(N)=\exp\!\Big(-N\exp\!\Big[\frac{2J_{0}^{2}}{1+2J_{0}}(N-1)-2h\Big]\Big),$ (66)
which provides a good approximation over a broad range of $J_{0}$ for sufficiently large $h$.
Appendix C Proof of Theorem 1
Proof.
Start with the condition
$N\geq N_{\mathrm{DC}}\;=\;1+\frac{1}{\log\alpha}\left(\log\left(1-\frac{2\log\theta_{N,k}}{\beta_{0}}\right)+\frac{\log 2}{1-1/k}\right),$ (67)
with $\alpha>1$, $\beta_{0}>0$, $0<\theta_{N,k}\leq 1$, $k\geq 2$. Each term in $N_{\mathrm{DC}}$ is non-negative,
$\log\alpha>0,\quad\log\left(1-\frac{2\log\theta_{N,k}}{\beta_{0}}\right)\geq 0,\quad\frac{\log 2}{1-1/k}>0,$ (68)
therefore Eq. (67) implies the following weaker conditions:
$N>1,$ (69)
$N>1+\frac{1}{\log\alpha}\log\left(1-\frac{2\log\theta_{N,k}}{\beta_{0}}\right)>1+\frac{1}{\log\alpha}\log\left(-\frac{2\log\theta_{N,k}}{\beta_{0}}\right),$ (70)
$N>\frac{1}{\log\alpha}\frac{\log 2}{1-1/k},$ (71)
which are nevertheless more directly useful.
Eq. (71) implies
$\begin{split}&N(1-1/k)\log\alpha>\log 2\\ \Rightarrow\;&\alpha^{N(1-1/k)}>2\\ \Rightarrow\;&\alpha^{-N(1-1/k)}<1/2\\ \Rightarrow\;&(1-\alpha^{N/k-N})>1/2.\end{split}$ (72)
Eq. (70) implies
$(N-1)\log\alpha>\log\left(-\frac{2\log\theta_{N,k}}{\beta_{0}}\right)\Rightarrow\;\frac{1}{2}\beta_{0}\alpha^{N-1}>-\log\theta_{N,k}.$ (73)
Putting Eqs. (69), (72), (73) together, we have
$\begin{split}\beta_{0}N(\alpha^{N-1}-\alpha^{N/k-1})&=\beta_{0}\alpha^{N-1}N(1-\alpha^{N/k-N})\\ &>\beta_{0}\alpha^{N-1}\times 1\times\frac{1}{2}\\ &>-\log\theta_{N,k},\end{split}$ (74)
therefore the logarithmic gain $\Delta(N,k)=\log\theta_{N,k}+\beta_{0}N(\alpha^{N-1}-\alpha^{N/k-1})>0$ is positive. ∎